It takes 9.5 hours to fly from Baltimore to Munich and 11 hours to return. If an airplane averages 550 mph in still air, what is the average rate of the wind blowing in the direction from Baltimore to Munich? Assume the speed of the wind (jet stream) is constant and the same for both legs of the trip.
Answer by rmnavalta(7)

This problem is a UNIFORM MOTION PROBLEM

We will be using distance= rate x time.

*Legend  B-Baltimore  M-Munich

Let x = average rate of wind blowing in the direction from B to M
Given: average rate of the airplane in still air = 550mph

Notes: With the wind, the rate of the airplane is equal to (550+x)mph. 
          -The plane moves faster.
       Against the wind, the rate of the airplane is equal to (550-x) mph.
          -The plane moves slower.

Table: 
         |  time(h)  |  rate(mph)  |   distance(mi)= rate*time   | 
B to M   |    9.5    |   550+x     |          9.5(550+x)         |
M to B   |    11     |   550-x     |           11(550-x)         |


       (Distance from B to M is the same as from M to B)

SET-UP THE EQUATION-->  9.5(550+x)=11(550-x)

Solve.  5225+9.5x       = 6050-11x
        5225+9.5x+11x   = 6050-11x + 11x   ---> Add 11x on both sides.
        5225+20.5x-5225 = 6050-5225        ---> Subtract 5225 on both sides.
          (20.5x)/20.5  = 825/20.5        ---> Divide both sides by 20.5.
                     x  = 40.24 mph

Answer: The average rate of wind blowing in the direction from B to M is approximately 40.24 mph.


God bless you! - email me @ rmnavalta@yahoo.com